Compute the mass defects of (a) 32S; (b) 20F; (c) 238U. 


1
Expert's answer
2021-06-14T17:26:59-0400

a)


"\Delta m=((16)1.00727+(32-16)1.00865)\\-31.972071=0.28265\ amu"

b)


"\Delta m=((9)1.00727+(20-9)1.00865)\\-19.99998=0.1606\ amu"

c)


"\Delta m=((92)1.00727+(238-92)1.00865)\\-238.05079=1.88095\ amu"